Phantomdex:

A large company tracks the transportation habits of its 300 employees. Their daily commuting distance and usual method are summarized in the table. Under 2 miles 2–4 miles 5–6 miles Over 6 miles Car 15 25 30 40 Public Transit 20 20 35 15 Bike 30 15 10 5 Walk 25 10 5 0 Part A: Calculate the empirical conditional probability of an employee who walks to work, given that they live under 2 miles away. Show all work. (2 points) Part B: Is an employee more or less likely to walk to work if they live under 2 miles away? Justify the answer mathematically. (2 points)

Phantomdex:

@eiwoh2 could I maybe get some help, none of the people I normally go to are on TwT

Phantomdex:

@midnight97

Phantomdex:

I don't know if you'll know how to do it, but toga isn't on so TwT

Midnight97:

Sorry Idk this one love

Phantomdex:

No one else is on TwT

Midnight97:

@aubree

Aubree:

One min

Aubree:

The empirical conditional probability of an employee walking to work when living under 2 miles away is approximately 27.78%. Employees are more likely to walk to work if they live under 2 miles away, as this probability is higher than the overall walking probability of approximately 13.33%.

Aubree:

To calculate the empirical conditional probability of an employee who walks to work, given that they live under 2 miles away, we divide the number of employees who walk and live under 2 miles by the total number of employees who live under 2 miles. In this case, there are 25 employees who walk to work and live under 2 miles away, and the total number of employees living under 2 miles away is 15 (Car) + 20 (Public Transit) + 30 (Bike) + 25 (Walk) = 90 employees. Therefore, the probability is calculated as follows: 25/90 = 0.2778 or approximately 27.78%. For part B, to determine if an employee is more or less likely to walk to work if they live under 2 miles away, we can compare the probability calculated in part A to the overall probability of walking to work for all distance categories. The total number of walkers is 25 (Under 2 miles) + 10 (2-4 miles) + 5 (5-6 miles) + 0 (Over 6 miles) = 40. The total number of employees is 300. So, the overall walking probability is 40/300 = 0.1333 or approximately 13.33%. Since 27.78% > 13.33%, we conclude that an employee is more likely to walk to work if they live under 2 miles away.

Phantomdex:

Thank you aubree, I keep wanting to call you cherry for some reason because your pfp from far off looks like cherri bomb TwT

Phantomdex:

( I know it's todoroki)

Aubree:

Mhm

Phantomdex:

But still, thank you ^^

Aubree:

Np